# Tomasz Gradowski Teodor Buchner

### v1

@@ 4 Discussion
First, let us discuss the results of the single-lead models. In these models, we take only one limb lead as input and try to reconstruct all the precordial leads. The quality of the reconstruction can be interpreted as the measure of the correlation between the input lead and the precordial leads. We must remember that all precordial leads have the same reference point, the Wilson central terminal, so the information from all the limb electrodes is present in all precordial leads to some extent. However, by comparing the results of the single-lead models, we can try to estimate the correlation between the single electrodes as well.
@@
The results show that leads V1 and V6 are reconstructed with the highest accuracy in all three scenarios (Fig. 2, first three rows). In particular, lead V6 appears to be more closely correlated with limb leads II and I. A slightly better reconstruction of V6 from lead II than from lead I suggests a stronger correlation between the potential on the V6 electrode and the potential on the LL electrode (left leg) than LA (left arm), since the electrode potential RA (right arm) is present in both leads I and II.
@@
The V3 and V4 leads were reconstructed with the lowest accuracy in all three scenarios (V4 has the lowest PCC value, V3 has the highest MSE value). This suggests that the potential on the V3 and V4 electrodes carries information that is not present in the limb leads. This is consistent with the fact that V3 is placed closest to the interventricular septum and is, on average, equally susceptible to potentials from both ventricles. Since it is centrally located on the chest, it captures a unique combination of signals that are not as prominent on the limb leads. The significance of the V3 lead will be seen in the 3-lead models.
To estimate the correlation between precordial leads, we trained models with two limb leads and one precordial lead as input (I + II + Vx). We also trained a model with two limb leads (I+II) as input, which serves as a baseline for the comparison.
The first observation is that precordial leads are the most correlated with the closest precordial leads, e.g., including the V2 lead in the input leads results in significantly better reconstruction of the V1 and V3 leads. This is very intuitive since the precordial electrodes are placed in a line along the heart’s axis, and the potential of one electrode is expected to be the most correlated with the potentials of the adjacent electrodes.
@@
The results show that the reconstruction of the precordial leads is the best when the V3 lead is used as input. This confirms the previous observation that the V3 lead carries the most information about the heart’s electrical activity, which is not present in the limb leads. The worst reconstruction is obtained when the V6 lead is used as an input. This observation is consistent with the previous one that the V6 lead is the most correlated with the limb leads and carries the least additional information. Including the V6 lead in the input leads does not affect the reconstruction of the V1-V4 leads and improves the reconstruction of only the V5 lead, measured from the closest electrode to the V6 electrode.
@@
The following models assume a scenario where we want to reconstruct all the precordial leads from two limb leads and two precordial leads. This scenario is possible in a clinical setting, where precordial leads are unavailable due to the patient’s condition or technical limitations. A performant model in this scenario would open the way for the development of a portable ECG device with fewer leads, which could be used in emergencies or remote areas. The decision on which ECG leads to include in such a device would be based on the correlation between the leads and the quality of the reconstruction. The PCC analysis gives us very close results with no decisive winner in this category. We show the combination I + II + V2 + V4 in Fig. 8, as it got the lowest MSE value. Leads V2 and V4 are positioned over areas central to the electrical activity of the heart. V2 is placed in the fourth intercostal space at the left sternal border, while V4 is placed in the fifth intercostal space at the midclavicular line. The proximity of these leads to critical cardiac structures allows them to capture significant electrical activity from both the anterior and septal regions of the heart. V2 and V4 provide a balanced view of the heart’s electrical activity; V2 captures data from the right ventricle and the septal area, while V4 captures signals from the anterior wall of the left ventricle. This complementary information is crucial for accurately reconstructing the missing leads. Although adjacent leads (like V2 and V3) provide highly correlated data, they can introduce redundancy rather than additional useful information. V2 and V4, being slightly more spatially separated, offer diverse and non-redundant information that enhances reconstruction accuracy.
@@ 5 Conclusion
The results of this study are promising and show the potential of deep learning models for the reconstruction of ECG. The method for estimating the amount of information carried by each lead and the correlations between the leads may be significant from the clinical point of view.
First, the proposed method can optimize ECG use. By quantifying the information carried by each lead, clinicians can prioritize specific leads over others. This can be especially useful when a full 12-lead ECG is impractical, such as in remote monitoring or emergency situations. Streamlining lead selection could improve diagnostic efficiency without compromising accuracy.
Second, the method can reduce redundancy. Correlations between leads can reveal which ones provide redundant information. By identifying this redundancy, clinicians can reduce the number of leads used in specific scenarios, potentially simplifying the equipment, reducing patient discomfort, and shortening ECG reading times.
Third, the method can improve personalized diagnostics. Different patients may have varying degrees of information spread across ECG leads. The presented method can help tailor the interpretation of the ECG to individual patients. It can also help detect subtle cardiac abnormalities.
Another value of the method is that it can provide a better understanding of cardiac physiology. Understanding the correlations between leads can provide deeper insight into the electrical activity of the heart and how it propagates across different axes, potentially aiding in a more accurate localization of ischemic or arrhythmic areas.
Finally, the method can improve telemedicine. For remote health monitoring, using fewer leads without sacrificing the quality of information could be a game changer. It could make ECG monitoring more accessible and affordable, especially in low-resource settings.
These advances can lead to faster diagnosis, more targeted treatment, and potentially lower healthcare costs.

### v2

@@ 4 Discussion
First, let us discuss the results of the single-lead models. In these models, we take only one limb lead as input and try to reconstruct all the precordial leads. The reconstruction quality can be interpreted as a measure of the correlation between the input and precordial leads. We must remember that all precordial leads have the same reference point, the Wilson central terminal, so the information from all the limb electrodes is present in all precordial leads to some extent. However, by comparing the single-lead models’ results, we can also try to estimate the correlation between the single electrodes.
@@
The results show that leads V1 and V6 are reconstructed with the highest accuracy in all three scenarios (Fig. 2, first three rows). In particular, lead V6 appears more closely correlated with limb leads II and I. A slightly better reconstruction of V6 from lead II than from lead I suggests a stronger correlation between the potential on the V6 electrode and the potential on the LL electrode (left leg) than LA (left arm), since the electrode potential RA (right arm) is present in both leads I and II.
@@
The V3 lead exhibited the lowest reconstruction accuracy in all three scenarios. This suggests that the potential on the V3 electrode carries information that is not present in the limb leads. This is consistent with that V3 is placed closest to the interventricular septum and is, on average, equally susceptible to potentials from both ventricles. Since it is centrally located on the chest, it captures a unique combination of signals not as prominent on the limb leads. The significance of the V3 lead will be seen in the 3-lead models.
To estimate the correlation between precordial leads, we trained models with two limb leads and one precordial lead as input (I + II + Vx). We also trained a model with two limb leads (I+II) as input, which serves as a baseline for the comparison. The first observation is that precordial leads are the most correlated with the closest precordial leads, e.g., including the V2 lead in the input leads results in significantly better reconstruction of the V1 and V3 leads. This is very intuitive since the precordial electrodes are placed in a line along the heart’s axis, and the potential of one electrode is expected to be the most correlated with the potentials of the adjacent electrodes. The results show that the reconstruction of the precordial leads is the best when the V3 lead is used as input. This confirms the previous observation that the V3 lead carries the most information about the heart’s electrical activity, which is not present in the limb leads. This observation aligns with the anatomical placement of V3 near the interventricular septum, which allows it to sample electrical activity from both ventricles. From a signal-processing perspective, its inclusion likely contributes orthogonal information that complements the limb leads. This reinforces the idea that lead configuration impacts the informational redundancy within ECG recordings, and has practical implications for lead selection in portable systems.
@@
The worst reconstruction is obtained when the V6 lead is used as an input. This observation is consistent with the previous one, which states that the V6 lead is the most correlated with the limb leads and carries the least additional information. Including the V6 lead in the input leads does not affect the reconstruction of the V1-V4 leads and improves the reconstruction of only the V5 lead, measured from the closest electrode to the V6 electrode.
@@
The following models assume a scenario where we want to reconstruct all the precordial leads from two limb leads and two precordial leads. This scenario is possible in a clinical setting, where precordial leads are unavailable due to the patient’s condition or technical limitations. A performant model in this scenario would open the way for developing a portable ECG device with fewer leads, which could be used in emergencies or remote areas. The decision on which ECG leads to include in such a device would be based on the correlation between the leads and the reconstruction quality. A naive approach would suggest that the best combination of leads should include lead V3, as it gave the best result in the previous models. However, the best combination of leads is I + II + V2 + V4. Leads V2 and V4 are positioned over areas central to the heart’s electrical activity. V2 is placed in the fourth intercostal space at the left sternal border, while V4 is placed in the fifth intercostal space at the midclavicular line. The proximity of these leads to critical cardiac structures allows them to capture significant electrical activity from both the anterior and septal regions of the heart. V2 and V4 provide a balanced view of the heart’s electrical activity; V2 captures data from the right ventricle and the septal area, while V4 captures signals from the anterior wall of the left ventricle. This complementary information is crucial for accurately reconstructing the missing leads. Although adjacent leads (like V2 and V3) provide highly correlated data, they can introduce redundancy rather than additional helpful information. V2 and V4, being slightly more spatially separated, offer diverse and non-redundant information that enhances reconstruction accuracy.
From a systems integration perspective, the model is well-suited for real-time or point-of-care deployment. The U-net architecture, as implemented in this study, maintains moderate parameter count and computational complexity. The fixed-length input segments are small enough to allow rapid inference. These characteristics make the proposed model highly compatible with bedside monitors, portable ECG systems, and remote health monitoring devices. Future work may involve prototyping deployments on representative hardware platforms to benchmark inference speed and memory usage formally. One limitation of the current study is that the model was trained and evaluated exclusively on clinical-grade ECG recordings obtained from standard 12-lead systems (PTB-XL and PTB datasets). Wearable ECG devices, such as smartwatches or portable monitors, typically exhibit different noise profiles, including motion artifacts, lower signal-to-noise ratios, and electrode placement variability. As a result, direct generalization to wearable device signals is not guaranteed. To achieve optimal generalization, future work could incorporate domain adaptation strategies, such as fine-tuning on wearable device datasets or augmenting training data with synthetic noise, to explicitly model the noise characteristics of wearable signals. This would enhance the model’s applicability to real-world telemedicine and mobile health scenarios.
@@ 5 Conclusion
The results of this study are promising and show the potential of deep learning models for the reconstruction of ECG. The method for estimating the amount of information carried by each lead and the correlations between the leads may be significant from the clinical point of view. First, the proposed method can optimize ECG use. By quantifying the information carried by each lead, clinicians can prioritize specific leads over others. This can be especially useful when a full 12-lead ECG, such as in remote monitoring or emergencies, is impractical. Streamlining lead selection could improve diagnostic efficiency without compromising accuracy. Second, the method can reduce redundancy. Correlations between leads can reveal which ones provide redundant information. By identifying this redundancy, clinicians can reduce the number of leads used in specific scenarios, potentially simplifying the equipment, reducing patient discomfort, and shortening ECG reading times. Third, the method can improve personalized diagnostics. Different patients may have varying degrees of information spread across ECG leads. The presented method can help tailor the interpretation of the ECG to individual patients. It can also help detect subtle cardiac abnormalities. Another value of the method is that it can provide a better understanding of cardiac physiology. Understanding the correlations between leads can offer more profound insight into the heart’s electrical activity and how it propagates across different axes, potentially aiding in a more accurate localization of ischemic or arrhythmic areas. Finally, the method can improve telemedicine. For remote health monitoring, using fewer leads without sacrificing the quality of information could be a game-changer. It could make ECG monitoring more accessible and affordable, especially in low-resource settings. These advances can lead to faster diagnosis, more targeted treatment, and potentially lower healthcare costs [27].